# Chemical strategies to mitigate electrostatic charging during coffee grinding

### 05 Dec 2023 v1

@@ 4 Aggregate formation
For the ionizing strategies, charge mitigation occurs at the chute exit, not in the grinder itself. Conversely, the water addition addresses charging throughout the process. These differences may have important effects on sheeting (i.e. wall adhesion), material loss, and clump formation, especially since previous work has shown that particles adhered to the wall can have Q/m ratios several dozen times larger than those forming the bulk.zhang2021experimental
Whether a particle of radius with charge will electrostatically adhere to a surface depends in part on its electrostatic-to-gravitational force ratio,
@@ 5 The effect of charge mitigation on espresso quality
Conversely, charge reduction using an ionizing source at the grinder chute does not produce a reduction in the mean particle size, nor does it decrease grinder retention (see Figures 4c and d). Thus, it is worth considering whether ionizer-based techniques produce departures from the nominal brewing behavior similar to those observed with added water. As evident in Figure 5b, we observe nearly indistinguishable differences in espresso shots prepared with or without the ionizers. The modest increase in %TDS from 7.76% to 8.02% (average over five replicates) is still notable, but the impact is far less pronounced. These data suggests that ionization techniques perhaps do cause some aggregates to break up as these exit the chute, but the effect is limited. However, they do not propitiate the reincorporation of small particles bound electrostatically to the interior surfaces of the grinder and burrs. Thus, unlike the water drop technique, discharging coffee at the nozzle produces limited changes to the physical characteristics of the grounds used to brew coffee compared to the untreated samples.
@@ 6 Conclusion
We have comprehensively assessed the performance of a broad range of electrostatic reduction techniques in the context of coffee grinding. Added moisture and high-voltage ionization effectively counteract charging generated through fracto- and triboelectric charging. At minimum, both techniques have the potential to decrease the gravimetric charge by at least 50 %. While ionization methods can reduce the Q/m ratios of expelled coffee to near zero if appropriately tuned to a given coffee, they generally do not address electrification processes and adhesion dynamics within the grinder. As such, ionization methods at the grinder chute do not mitigate material loss (retention) which, for a grinder like the EK43, can exceed 10%. Conversely, the charge reduction afforded by the water addition (involving volumes ranging from 0-50 L g-1 dry mass coffee) has the capacity to resolve aggregation effects across the entire grinding system, even if particles retain some charge. Because particle-wall aggregates comprise smaller grains, we find that reincorporating this material into the bulk significantly changes the espresso brewing behavior (slower shots and smaller flow rates) and resultant increase in coffee solubility. And while our experiments conclusively demonstrate the economic upside of water incorporation during grinding for espresso, we suspect a similar effect will be observed for all percolation brewing techniques, as liberation of fine particles will not only clog void space in the coffee bed, but also fill voids in filter paper in pour-over embodiments.
@@ Acknowledgements
JMH acknowledges support from DOE Grant No. 242649. CHH acknowledges the support from the National Science Foundation under Grant No. 2237345 and support from the Camille and Henry Dreyfus Foundation. This work was supported by the Coffee Science Foundation, underwritten by Nuova Simonelli. The work was enabled by donations of green coffee from Farmers Union and Finca La Ilusión. We are grateful to Ikawa for providing the roaster, Pentair for providing the reverse osmosis water filtration system, and Tailored Coffee, Eugene, OR for the use of the EK43.

### 19 Dec 2023 v3

@@ 4 Aggregate formation
For ionizing strategies, charge mitigation occurs at the chute exit, not in the grinder itself. Conversely, the water addition evidently passivates charging throughout the grinding process. These differences may have important effects on particle-wall adhesion, material loss, and clump formation, especially since previous work has shown that particles adhered to the wall can have Q/m ratios several dozen times larger than those forming the bulk.zhang2021experimental
Whether a particle of radius with charge will electrostatically adhere to a surface depends in part on its electrostatic-to-gravitational force ratio (EGR),
@@ 5 The effect of charge mitigation on espresso quality
Conversely, charge reduction using an ionizing source at the grinder chute does not produce a reduction in the mean particle size, nor does it decrease grinder retention (see Figures 4c and d). As evident in Figure 5b, we observe nearly indistinguishable differences in espresso shots prepared with or without the ionizers. The modest increase in %TDS from 7.76% to 8.02% (average over five replicates) is still notable, and can be attributed to some fine particles being liberated from boulders. These data suggests that ionization techniques perhaps do cause some aggregates to break up as these exit the chute, but the effect is limited. However, they do not propitiate the reincorporation of small particles bound electrostatically to the interior surfaces of the grinder and burrs. Thus, unlike the addition of water, discharging coffee at the nozzle produces limited changes to the physical characteristics of the grounds used to brew coffee compared to the untreated samples.
@@ 6 Conclusion
We have assessed the performance of a range of electrostatic reduction techniques in the context of coffee grinding. Added moisture and high-voltage ionization effectively counteract charging generated through fracto- and triboelectric charging. At minimum, both techniques have the potential to decrease the gravimetric charge by at least 50 %. Unipolar ionization methods can reduce the Q/m ratios of expelled coffee to near zero if appropriately tuned to a given coffee. Bipolar or balanced ionization may be less effective than unipolar ionization, but can produce charge reductions comparable to those afforded by water addition. However, ionization generally does not address electrification processes and adhesion dynamics within the grinder. As such, ionization methods at the grinder chute do not mitigate material loss (retention) which, for a grinder like the EK43, can be hundreds of milligrams of dry material. Conversely, the charge reduction afforded by water addition (volumes ranging from 0-50 L g-1 dry mass coffee) has the capacity to resolve aggregation effects across the entire grinding system, even if particles retain some charge. Because particle-wall aggregates comprise smaller grains, we find that reincorporating this material into the bulk significantly changes the espresso brewing behavior (slower shots and smaller flow rates) and resultant increase in coffee solubility. And while our experiments conclusively demonstrate the economic upside of water incorporation during grinding for espresso, we suspect a similar effect will be observed for all percolation brewing techniques, as liberation of fine particles will not only clog void space in the coffee bed, but also fill voids in filter paper in pour-over embodiments.
@@ Acknowledgements
JMH acknowledges support from start-up funds at Portland State University. CHH acknowledges the support from the National Science Foundation under Grant No. 2237345 and support from the Camille and Henry Dreyfus Foundation. This work was supported by the Coffee Science Foundation, underwritten by Nuova Simonelli. The work was enabled by donations of green coffee from Farmers Union and Finca La Ilusión. We are grateful to Ikawa for providing the roaster, Pentair for providing the reverse osmosis water filtration system, and Tailored Coffee, Eugene, OR for the use of the EK43.